# persiansort : an alternative to mergesort inspired by persian rug

### v2

@@ Abstract
This paper introduces persiansort, new stable sorting algorithm inspired by Persian rug. Persiansort does not have the weaknesses of mergesort under scenarios involving nearly sorted and partially sorted data, also utilizing less auxiliary memory than mergesort and take advantage of runs. Initial experimental showed, this method is flexible, powerful and works better than mergesort in almost all types of data. Persiansort offers several advantages over merge methods, make it a potential replacement.
@@ 3 New Algorithm Design
The initial phase in developing the new algorithm is the division or warping stage. We select the top-down warp configuration for this purpose. In the top-down mergesort technique, the data is bifurcated at each stage but we do this with more flexibility so can make more effective knots. To achieve this objective, we divide the data into further segments (”wp” segments). The number of segments (wp) can even change in different conditions and parts of the algorithm. We obtain the size of each section from the following function: $((end-start+1)/wp)+1$
In persiansort, warping (divide step) can transition from wp=4 to wp=n+1 where ”n” is the size of the original dataset. Alternative functions for warping may also be utilized, for instance: max-jump= $\sqrt{(end-start+1)}+1$.
@@ 4 Experimental
To evaluate the effectiveness of persiansort and juxtapose this approach with other methods, we designated the wp value as 9, acknowledging that the optimal wp may vary and superior results may be achieved with different amount of wp. To evaluate the efficacy of this method and juxtapose it with alternative methods, we employ the same technique utilized for eqsort and sort a substantial quantity of arrays, analyzing the average time for varying data sizes [1]. The compared methods are selected as follows:
1.Persiansort method utilizing the original algorithm without hybridization, with wp=9. In persiansort for K-Nearly Sorted, we selected $wp=n/(1.5\sqrt{k})$ and hybridized it with insertion sort. In partial sorting, we employed the ideal wp.
@@ 2. Mergesort2 (M2) with n/2 auxiliary memory:
mergesort(int ar[],int aux[] ,int start,int end){
if(start>=end) return;
if(check_descending==true) {reverse(ar,start,end); return;}
int m=(start+end)/2;
mergesort(ar,aux,start,m)
mergesort(ar,aux,m+1,end)
if(ar[m+1]>=ar[m]) return;
merge(ar,aux,start,m,end);
3.Mergesort1 (M1) with n auxiliary memory with the subsequent algorithm.
mergesort(int a[],int b[],int start,int end){
if(start>=end) return;
if(check_descending==true) {reverse(a,start,end); return;}
int m=(start+end)/2;
mergesort(b,a,start,m)
mergesort(b,a,m+1,end)
if(b[m+1]>=b[m]) {coppybtoa(a,b,start,end); return;}
merge(a,b,start,m,end);
@@ 5 Random Data
The findings for random data indicate that the M1 algorithm is a little faster than other algorithm. M1 uses additional memory as O(n) and other results show that this method does not perform well in other conditions. Timsort is a hybrid, derived from merge sort and insertion sort. The novel persiansort approach can fully compete with Merge methods in random data at the optimal wp value and the ideal hybrid value for small datasets, requiring significantly less auxiliary memory (n/wp).
@@ 6 Data with runs
To evaluate the efficacy of the novel algorithm against several methods, we generate datasets including one million entries, ensuring that 60 percent of this data contains runs of size R (in ascending or descending order). For each measurement, we organize 1000 distinct sets under the stated parameters for each approach and calculate the average sorting time for each technique.
3030303050505050100100100100200200200200500500500500100010001000100020002000200020005000500050005000100001000010000100000.70.70.70.70.80.80.80.80.90.90.90.911111.11.11.11.11.21.21.21.21.31.31.31.3R (size of ascending runs)Runtime ratioT-M2 / T-M2T-M1 / T-M2T-TIMS / T-M2T-PER / T-M2
3030303050505050100100100100200200200200500500500500100010001000100020002000200020005000500050005000100001000010000100000.70.70.70.70.80.80.80.80.90.90.90.911111.11.11.11.11.21.21.21.21.31.31.31.3R (size of descending runs)T-M2 / T-M2T-M1 / T-M2T-TIMS / T-M2T-PER / T-M2
@@
Figure 6 illustrates that the effectiveness of the M1 technique declines significantly when the sort percentage of the dataset or the size of the runs increases, rendering this method an unsuitable option for this type of data. The remaining three approaches exhibit comparable speeds and efficiency. The persiansort approach has somewhat better performance as the sorting percentage and the size of the runs rise. Figure 9 displays the results for 30 and 90 percent.
@@ 7 Nearly Sorted data
We assess nearly sorted data in two sections to analyze and compare the persiansort approach with other techniques. In K-Nearly, each element is at most k positions away from its correct sorted position. In this kind of data, for persiansort, we designate the wp value as $n/(1.5\sqrt{k})$. In addition to the methodologies from prior stages, we also examine the insertion sort technique for K-Nearly. The data in Figure 7b indicates that the persiansort approach significantly outperforms merge’s methods. An analysis of the two rapid methodologies reveals that:
6570758085909396990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.4 pp (percent of pre-sorted)Runtime ratioT-M2 / T-M2T-M1 / T-M2T-TIMS / T-M2T-PER / T-M2
2510203550751002000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.41.61.61.61.6 k T-M2 / T-M2T-M1 / T-M2T-TIMS / T-M2T-INS / T-M2T-PER / T-M2
Figure 7: (a): Runtime ratio for dataset of size $10^{6}$ for increasing percent of pre-sorted data. (b): Runtime ratio for incresing k for k-nearly sorted data with 5 methods.
  • The persiansort algorithm, akin to insertion sort, uses minimal auxiliary memory. persiansort utilizes memory equivalent to $31.5\sqrt{k}+$ 3.
@@
A distinct kind of nearly sorted data is that in which the initial segment is ordered, exemplified by the addition of new data to an already sorted dataset, like online algorithms. We evaluated the novel persiansort approach against merging methods across various percentages of pre-sorted data. Figure 7a illustrates that as the amount of pre-sorted data increases, persiansort significantly outperforms other approaches, achieving nearly double the speed of mergesort methods at large pre-sorted percentages.
8 partial sorting
Partial sorting is returning a list of the k smallest (or k largest) elements in order. We analyzed and contrasted the new persiansort approach with the merge methods in partial sorting. In this iteration of persiansort, we selected the wp value based on the table below and implemented minor modifications across all comparable algorithms Suitable for this type of data.
k 100 90 80 70 60 50 40 30 20 10 wp n /16.7 n /16.1 n /15.4 n /14.3 n /13.3 n /11.7 n /10.5 n /8.3 n /5 n /3.3
The results indicate that the new method exhibits significantly enhanced efficiency, particularly for smaller values necessitated for sorting.
1009080706050403020100.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.911111.11.11.11.11.21.21.21.21.31.31.31.3k (number of element for every part)Runtime ratioT-M2 / T-M2T-TIMS / T-M2T-PER / T-M2
Figure 8: Runtime ratio for number of element must be sorted by three methods.

### v3

@@ Abstract
This paper introduces persiansort, new stable sorting algorithm inspired by Persian rug. Persiansort does not have the weaknesses of mergesort under scenarios involving nearly sorted and partially sorted data, also utilizing less auxiliary memory than mergesort and take advantage of runs. Initial experimental showed, this method is flexible, powerful and works better than mergesort in almost all types of data. Persiansort offers several advantages over merge methods make it a potential replacement.
@@ 3 New Algorithm Design
The initial phase in developing the new algorithm is the division or warping stage. We select the top-down warp configuration for this purpose. In the top-down mergesort technique, the data is bifurcated at each stage but we do this with more flexibility so can make more effective knots. To achieve this objective, we divide the data into further segments (”wp” segments). The number of segments (wp) can even change in different conditions and parts of the algorithm. We obtain the size of each section from the following function: ((endstart+1)/wp)+1((end-start+1)/wp)+1 .
In persiansort, warping (divide step) can transition from wp=4 to wp=n+1 where ”n” is the size of the original dataset. Alternative functions for warping may also be utilized.
@@
4 Time and Space complexity
Due to the presence of the ascending and descending knots in lines 22 and 25 of the algorithm, the best-case time complexity of the algorithm is O(n), which occurs when the input data is already sorted. Furthermore, when the data is nearly sorted, the algorithm maintains a linear time complexity of O(n). Conversely, the worst-case time complexity arises when the data is completely random, such that the performance of knots is at its minimum. In this situation, the worst-case time complexity increases to O(nlogn).
The amount of memory required for this sorting method entirely depends on the selected value of wp and is determined by the function auxiliary-size. The required memory is given by the equation: $n/wp+n/(wp)^{2}+n/(wp)^{3}+$... Nevertheless, if necessary, persiansort can be implemented in such a way that it requires only $n/wp$ .
This algorithm demonstrates considerable flexibility and robustness, as optimal conditions for sorting various types of data can be achieved by adjusting a single parameter, denoted as wp. From a theoretical standpoint, considering that for random datasets the number of comparisons in the persiansort algorithm can be approximately expressed by the relation $((wp)/2)n\log_{wp}n$ (assuming the impact of knots operations is neglected), the optimal value of wp is determined to be 4. However, due to the influence of knots, the actual optimal value of wp is expected to differ. Preliminary analyses and computations indicate that the optimal range for wp for the random dataset is estimated to be between 6 and 11(this value also depends on the parameter n). Selecting any value within this range results in a performance variation of only about 0–5 percent. Therefore, considering both performance stability and reduced memory consumption, the value 9 was chosen for wp.
For k-nearly sorted datasets, the parameter max-jump is typically set to a value approximately equal to k, and initial evaluations revealed that choosing an appropriate $wp=n/(1.5\sqrt{k})$ value enhances the efficiency of the algorithm.
@@ 5 Experimental
To evaluate the effectiveness of persiansort and juxtapose this approach with other methods, we designated the wp value as 9. To evaluate the efficacy of this method and juxtapose it with alternative methods, we employ the same technique utilized for eqsort and sort a substantial quantity of arrays, analyzing the average time for varying data sizes [9]. For the purpose of comparing the sorting times of persiansort with other techniques, a large number of datasets sorted. the data type has been set to double and the average sorting time is computed. The number of datasets is increased until a four-digit reproducibility is achieved, corresponding to an error of less than 0.1 percent.
All algorithms have been implemented in the C++ programming language and have been compiled with the GNU C++ Compiler 11.4.0 . All tests have been performed on a PC with an i3-380M , 2 core , 2.53 GHz, with 4 GB memory and running Linux Ubuntu 22.04.
The compared methods are selected as follows:
1.Persiansort method utilizing the original algorithm without hybridization, with wp=9.
In persiansort for K-Nearly Sorted, we selected $wp=n/(1.5\sqrt{k})$.
@@ 2. Mergesort2 (M2) with n/2 auxiliary memory:
3.Mergesort1 (M1) with n auxiliary memory.
@@ 6 Random Data
The findings for random data indicate that the M1 algorithm is a little faster than other algorithm. M1 uses additional memory as O(n) and other results show that this method does not perform well in other conditions. Timsort is a hybrid, derived from merge sort and insertion sort. The novel persiansort approach can fully compete with Merge methods in random data, requiring significantly less auxiliary memory (n/wp).
@@ 7 Data with runs
To evaluate the efficacy of the novel algorithm against several methods, we generate datasets including one million entries, ensuring that 60 percent of this data contains runs of size R (in ascending or descending order). For each measurement, we organize 1000 distinct sets under the stated parameters for each approach and calculate the average sorting time for each technique. Figure 6 and other result illustrates that the effectiveness of the M1 technique declines significantly when the sort percentage of the dataset or the size of the runs increases, rendering this method an unsuitable option for this type of data. The remaining three approaches exhibit comparable speeds and efficiency. The persiansort approach has somewhat better performance as the sorting percentage and the size of the runs rise.
3030505010010020020050050010001000200020005000500010000100000.70.70.80.80.90.9111.11.11.21.21.31.3R (size of ascending runs)Runtime ratioT-M2 / T-M2T-M1 / T-M2T-TIMS / T-M2T-PER / T-M2
3030505010010020020050050010001000200020005000500010000100000.70.70.80.80.90.9111.11.11.21.21.31.3R (size of descending runs)T-M2 / T-M2T-M1 / T-M2T-TIMS / T-M2T-PER / T-M2
@@ 8 Nearly Sorted data
We assess nearly sorted data in two sections to analyze and compare the persiansort approach with other techniques. In K-Nearly, each element is at most k positions away from its correct sorted position. In this kind of data, for persiansort, we designate the wp value as $n/(1.5\sqrt{k})$. In addition to the methodologies from prior stages, we also examine the insertion sort technique for K-Nearly.
K T-M1 T-M2 T-TIMS T-INS T-PER
2 0.140 0.103 0.075 0.014 0.014
5 0.163 0.151 0.109 0.019 0.018
10 0.170 0.159 0.120 0.031 0.031
20 0.168 0.161 0.135 0.054 0.047
35 0.173 0.166 0.143 0.080 0.046
50 0.174 0.166 0.144 0.109 0.062
75 0.173 0.167 0.146 0.153 0.068
100 0.173 0.169 0.146 0.196 0.073
200 0.173 0.170 0.148 0.370 0.089
Table 1: Runtime for increasing k for k-nearly sorted data.
2510203550751002000.20.20.40.40.60.60.80.8111.21.21.41.41.61.6 k RatioT-M2 / T-M2T-M1 / T-M2T-TIMS / T-M2T-INS / T-M2T-PER / T-M2
Figure 7: Runtime ratio
The results in Figure 7 indicate that the persiansort approach significantly outperforms merge’s methods. An analysis of the two rapid methodologies reveals that:
  • The persiansort algorithm, akin to insertion sort, uses minimal auxiliary memory. persiansort utilizes memory equivalent to $1.5\sqrt{k}+3$.
@@
A distinct kind of nearly sorted data is that in which the initial segment is ordered, exemplified by the addition of new data to an already sorted dataset, like online algorithms. We evaluated the novel persiansort approach against merging methods across various percentages of pre-sorted data.
[…]
65 0.168 0.122 0.113 0.118
70 0.160 0.109 0.099 0.100
75 0.151 0.093 0.087 0.087
80 0.143 0.081 0.072 0.069
85 0.135 0.069 0.060 0.055
90 0.127 0.056 0.046 0.037
93 0.123 0.049 0.039 0.029
96 0.118 0.041 0.031 0.021
99 0.113 0.035 0.024 0.013
Table 2: Runtime for dataset of size $10^{6}$.
6570758085909396990.20.20.40.40.60.60.80.8111.21.21.41.4 $p$ (percent of pre-sorted)RatioT-M2 / T-M2T-M1 / T-M2T-TIMS / T-M2T-PER / T-M2
Figure 8: Runtime ratio
Figure 8 illustrates that as the amount of pre-sorted data increases, persiansort significantly outperforms other approaches, achieving nearly double the speed of mergesort methods at large pre-sorted percentages.